OpenStudy (anonymous):

If the net price is $900 and the trade discount rate is 40 percent, the list price would be:

OpenStudy (anonymous):

350

OpenStudy (anonymous):

I meant 375

OpenStudy (anonymous):

I think it should be $1500: \[\Large \frac{{900}}{{(100 - 40)/100}} = \frac{{900}}{{0.6}} = 1500\]